# A Closer Look on Memorization in Tabular Diffusion Model: A Data-Centric Perspective

### v1

@@ Abstract
Diffusion models have shown strong performance in generating high-quality tabular data, but they carry privacy risks by inadvertently reproducing exact training samples. While prior work focuses on dataset-level augmentation for memorization mitigation, little is known about which individual samples contribute most to memorization. In this paper, we present the first data-centric study of memorization dynamics in tabular diffusion models. We begin by quantifying memorization for each real sample based on how many generated samples are flagged as its memorized replicas, using a relative distance ratio metric.111We follow Yoon et al. (2023b); Gu et al. (2023); Fang et al. (2024) to define sample memorization using the relative distance ratio. Our empirical analysis reveals a heavy-tailed distribution of memorization counts: a small subset of samples disproportionately contributes to leakage, a finding further validated through sample-removal experiments. To better understand this effect, we divide real samples into the top- and non-top-memorized groups (tags) and analyze their training-time behavior differences. We track when each sample is first memorized and monitor per-epoch memorization intensity (AUC) across groups. We find that memorized samples tend to be memorized slightly earlier and show significantly stronger memorization signals in early training stages. Based on these insights, we propose DynamicCut, a two-stage, model-agnostic mitigation method. DynamicCut (a) ranks real samples by their epoch-wise memorization intensity, (b) prunes a tunable top fraction, and (c) retrains the model on the filtered dataset. Across multiple benchmark tabular datasets and tabular diffusion models, DynamicCut reduces memorization ratios with negligible impact on data diversity and downstream task performance, and complements existing data augmentation methods for further memorization mitigation. Furthermore, DynamicCut has transferability across different generative models for memorization sample tagging, i.e., high-ranked samples identified from one model (e.g., a diffusion model) are also effective in reducing memorization when removed from other generative models such as GANs and VAEs.
@@ 1 Introduction
Despite the impressive performance of diffusion models in synthesizing high-quality tabular data, memorization remains a critical concern. In particular, tabular generative models risk reproducing individual training records that, while appearing structurally altered, still retain sensitive attributes of real individuals. This subtle form of privacy leakage violates core principles of generative modeling and has severe implications in sensitive domains such as healthcare and finance Karras et al. (2022); Carlini et al. (2021); Song et al. (2021); Ho et al. (2020). To address this, prior work has primarily focused on memorization detection based on nearest-neighbor distance ratios Yoon et al. (2023a) —or on dataset-level augmentation strategies such as feature mixing TabCutMix Fang et al. (2024). However, these methods fall short in two key ways. First, the memorization detection tool does not explain when or which samples are memorized during training. Second, the memorization mitigation method like TabCutMix operates uniformly across the dataset, ignoring the heterogeneous nature of sample-level memorization risks. This gap in understanding motivates a central question:
@@ 2 Related Work
Memorization has been widely studied in image and language generation van den Burg and Williams (2021); Gu et al. (2023); Huang et al. (2024), where models often replicate training data. In vision, diffusion models like Stable Diffusion Rombach et al. (2022) and DDPM Ho et al. (2020) exhibit memorization Somepalli et al. (2023); Carlini et al. (2021), mitigated by methods such as concept ablation Kumari et al. (2023) and adaptive sample suppression Chen et al. (2024). In language models, techniques like Goldfish Loss Hans et al. (2024) and early memorization prediction Biderman et al. (2024) aim to reduce overfitting to specific sequences. However, memorization in tabular data generation remains underexplored. Due to the structured and mixed-type nature of tabular data, understanding how memorization arises—and how it can be mitigated—presents a distinct and open research challenge.
@@ 4.1 Memorization and Forgetting: Event Definitions
We adopt the widely used relative distance ratio criterion Yoon et al. (2023b); Gu et al. (2023); Fang et al. (2024) to determine whether a generated sample memorizes a real training sample. Formally, let be the training dataset and be a distance metric in the input space. Let and denote the nearest and second-nearest training samples to , respectively. We define the distance ratio as , A sample is considered memorized if , where is a fixed memorization threshold (we follow prior work and use ).
@@ 5.2 Algorithms
cell21 = r=6,
cell81 = r=6,
cell141 = r=6,
cell201 = r=6,
cell261 = r=6,
hline1-2,8,14,20,26 = -,
& Methods Mem. Ratio (%) Improv. MLE (%) -Precision(%) -Recall(%) Shape Score(%) Trend Score(%) C2ST(%) DCR(%)
Default
TabDDPM 0.45 - 76.79 0.69 98.15 1.45 44.41 0.70 97.58 0.95 94.46 0.68 91.85 6.04 49.12 0.94
TabDDPM+DC 1.04 76.42 0.73 93.26 5.44 42.64 1.16 93.40 3.46 89.65 4.00 86.09 6.48 50.78 1.60
CTGAN 0.63 - 68.68 0.21 68.95 1.70 16.49 0.55 85.04 0.93 77.05 2.47 61.89 4.89 49.45 0.80
CTGAN+DC 0.20 70.04 0.66 69.06 3.69 16.46 0.97 85.11 0.20 77.03 0.41 58.03 5.65 50.93 1.82
TVAE 0.43 - 72.24 0.36 82.97 0.37 20.57 0.42 89.37 0.54 83.36 0.99 52.63 0.37 51.33 0.96
TVAE+DC 0.60 76.87 0.54 82.54 1.97 21.25 0.60 89.98 0.74 80.36 0.63 51.42 3.67 50.01 2.21
Shoppers
TabDDPM 0.31 - 92.17 0.32 93.16 1.58 52.57 1.30 97.08 0.46 92.92 3.27 86.74 0.63 51.36 0.63
TabDDPM+DC 2.38 91.08 1.18 91.88 2.74 52.74 2.67 93.56 2.89 89.40 3.87 80.12 4.54 51.87 0.23
CTGAN 1.15 - 83.22 1.30 85.26 5.17 26.67 1.66 77.73 0.50 86.48 0.69 67.18 5.84 48.54 0.74
CTGAN+DC 0.45 83.71 0.44 89.13 4.17 27.73 0.66 77.78 0.87 86.93 0.34 69.31 6.11 48.55 1.50
TVAE 0.77 - 86.65 0.48 50.32 2.98 11.16 1.33 74.75 0.92 77.56 1.11 21.29 3.28 43.31 0.65
TVAE+DC 0.83 86.93 0.38 50.61 1.39 12.49 0.85 74.66 0.73 79.14 0.30 21.38 2.28 47.61 5.38
Adult
TabDDPM 0.18 - 91.09 0.07 93.58 1.99 51.52 2.29 98.84 0.03 97.78 0.07 94.63 1.19 51.56 0.34
TabDDPM+DC 1.03 90.81 0.29 97.45 1.20 45.54 1.32 95.68 2.55 93.62 2.82 86.84 3.48 50.42 1.24
CTGAN 0.62 - 88.64 0.32 78.15 3.66 26.27 0.67 82.43 0.90 82.83 0.93 63.64 2.74 49.14 0.27
CTGAN+DC 0.27 88.23 0.67 74.41 1.07 23.52 0.92 81.08 1.56 83.26 1.00 61.72 4.20 48.14 0.51
TVAE 0.41 - 88.61 0.49 92.34 2.19 29.90 1.00 82.71 0.45 79.06 0.90 48.67 2.70 48.76 0.25
TVAE+DC 1.10 88.25 0.33 87.80 0.74 30.57 2.82 83.54 0.56 78.71 0.86 55.64 2.59 45.96 0.54
@@ 6.2 Unified Evaluation of Memorization under Data Quality and Training Dynamics
cell21 = r=15,
cell171 = r=15,
cell321 = r=15,
cell471 = r=15,
cell621 = r=15,
hline1-2,17,32,47,62 = -,
@@
& Methods Mem. Ratio (%) Improv. MLE (%) -Precision(%) -Recall(%) Shape Score(%) Trend Score(%) C2ST(%) DCR(%)
Default
TabDDPM 0.45 - 76.79 0.69 98.15 1.45 44.41 0.70 97.58 0.95 94.46 0.68 91.85 6.04 49.12 0.94
TabDDPM+SMOTE 0.51 76.92 0.35 91.19 0.68 40.52 0.65 94.89 1.46 28.63 2.28 72.73 0.69 50.95 0.38
TabDDPM+TCM 0.47 76.47 0.60 97.30 0.46 38.72 2.78 97.27 1.74 93.27 2.52 94.72 3.87 50.23 0.53
TabDDPM+TCMP 0.24 76.92 0.17 98.26 0.25 41.92 0.52 97.37 0.09 91.42 1.15 95.64 0.49 49.75 0.32
TabDDPM+DCM 2.06 76.68 0.76 91.97 1.78 37.89 0.84 93.62 2.78 89.30 3.02 87.63 2.15 48.17 2.59
CTGAN 0.63 - 68.68 0.21 68.95 1.70 16.49 0.55 85.04 0.93 77.05 2.47 61.89 4.89 49.45 0.80
CTGAN+SMOTE 0.41 71.23 1.54 65.18 2.81 17.13 0.20 84.31 0.20 31.76 0.11 56.35 0.60 49.05 2.85
CTGAN+TCM 0.28 73.21 0.45 75.30 2.12 16.63 0.27 85.77 0.13 76.69 1.20 58.62 5.34 49.90 3.12
CTGAN+TCMP 0.25 70.05 0.81 71.33 0.82 17.02 0.60 85.28 0.97 78.09 0.32 63.67 3.07 50.14 0.60
CTGAN+DCM 0.07 72.37 0.39 71.72 0.74 16.64 0.20 85.78 0.70 76.85 1.62 57.98 3.09 48.48 0.24
TVAE 0.43 - 72.24 0.36 82.97 0.37 20.57 0.42 89.37 0.54 83.36 0.99 52.63 0.37 51.33 0.96
TVAE+SMOTE 0.41 75.41 0.42 82.47 0.42 20.77 0.73 88.79 0.08 79.34 2.40 53.28 3.26 49.20 1.54
TVAE+TCM 0.37 75.05 0.66 80.23 2.07 21.90 2.36 89.65 0.15 80.15 0.75 52.10 4.29 49.99 1.78
TVAE+TCMP 0.28 72.75 0.52 81.57 0.32 19.52 0.39 89.33 0.48 78.04 6.05 45.60 0.90 50.07 0.52
TVAE+DCM 0.21 76.27 0.25 85.49 0.41 21.66 0.19 90.36 0.25 86.76 5.13 54.81 0.79 51.07 1.35
Shoppers
TabDDPM 0.31 - 92.17 0.32 93.16 1.58 52.57 1.30 97.08 0.46 92.92 3.27 86.74 0.63 51.36 0.63
TabDDPM+SMOTE 1.46 89.96 0.95 94.41 4.67 45.22 3.26 90.78 0.49 83.09 2.47 64.05 1.44 51.94 1.52
TabDDPM+TCM 1.17 92.17 0.26 94.41 1.49 50.05 1.59 97.18 0.34 93.95 0.51 86.96 0.50 47.52 1.81
TabDDPM+TCMP 0.35 92.09 0.99 93.43 1.65 52.30 0.73 97.31 0.22 94.79 0.30 87.02 2.04 50.83 0.59
TabDDPM+DCM 2.83 91.46 0.74 93.25 3.87 46.15 4.42 92.32 5.00 89.11 4.95 80.56 4.81 49.74 3.34
CTGAN 1.15 - 83.22 1.30 85.26 5.17 26.67 1.66 77.73 0.50 86.48 0.69 67.18 5.84 48.54 0.74
CTGAN+SMOTE 1.86 87.09 1.22 76.13 13.12 26.19 3.66 75.57 2.48 83.94 1.13 59.83 6.67 50.74 1.75
CTGAN+TCM 0.42 84.21 1.21 80.82 7.68 26.50 1.50 77.51 2.15 85.75 1.31 62.04 6.75 49.90 2.33
CTGAN+TCMP 0.35 83.94 0.80 82.07 9.62 23.47 0.43 76.15 0.22 84.56 0.40 64.60 0.82 52.02 0.13
CTGAN+DCM 0.56 85.51 0.46 83.51 1.60 26.19 0.82 78.23 0.30 86.36 0.47 68.85 2.45 50.64 1.11
TVAE 0.77 - 86.65 0.48 50.32 2.98 11.16 1.33 74.75 0.92 77.56 1.11 21.29 3.28 43.31 0.65
TVAE+SMOTE 0.39 89.71 0.69 52.39 3.94 13.74 0.54 71.80 0.63 76.20 0.68 20.67 2.34 47.42 3.40
TVAE+TCM 2.38 87.33 0.90 49.75 5.04 12.66 3.18 75.47 1.30 78.91 1.30 21.87 3.87 50.06 1.78
TVAE+TCMP 0.80 87.07 0.74 50.69 0.44 10.54 1.68 74.86 0.26 78.55 0.48 21.78 4.28 43.21 0.69
TVAE+DCM 1.33 88.32 0.29 51.12 2.46 11.98 0.58 75.02 0.97 79.02 0.12 22.99 2.12 49.98 2.55
Adult
TabDDPM 0.18 - 91.09 0.07 93.58 1.99 51.52 2.29 98.84 0.03 97.78 0.07 94.63 1.19 51.56 0.34
TabDDPM+SMOTE 0.78 90.41 0.36 94.93 1.72 46.10 0.65 93.40 1.12 90.76 1.76 80.75 0.84 51.82 0.56
TabDDPM+TCM 0.19 91.15 0.06 94.97 0.06 47.43 1.46 98.65 0.03 97.75 0.07 85.61 16.03 50.99 0.65
TabDDPM+TCMP 2.11 90.54 0.17 92.26 6.97 43.49 3.74 95.10 4.27 91.50 6.53 84.76 10.12 50.68 0.89
TabDDPM+DCM 0.57 90.81 0.16 95.23 1.14 45.76 1.02 98.27 0.32 96.71 0.86 92.64 1.65 50.79 0.46
CTGAN 0.62 - 88.64 0.32 78.15 3.66 26.27 0.67 82.43 0.90 82.83 0.93 63.64 2.74 49.14 0.27
CTGAN+SMOTE 0.69 88.59 0.41 74.96 5.01 24.65 1.44 82.16 1.14 80.23 0.38 60.01 2.10 51.71 2.15
CTGAN+TCM 0.48 88.94 0.64 76.10 4.79 26.29 1.53 83.26 1.30 80.09 1.89 61.16 3.59 49.87 0.64
CTGAN+TCMP 0.31 88.63 0.66 76.27 0.60 25.56 0.33 81.42 0.32 82.11 1.01 60.98 1.67 50.96 0.17
CTGAN+DCM 0.88 88.45 0.29 79.31 1.04 23.31 1.42 82.23 0.23 82.80 2.27 63.46 3.43 48.13 0.33
TVAE 0.41 - 88.61 0.49 92.34 2.19 29.90 1.00 82.71 0.45 79.06 0.90 48.67 2.70 48.76 0.25
TVAE+SMOTE 0.16 87.89 0.32 89.00 3.48 31.01 2.50 80.65 1.28 81.83 1.64 52.98 6.01 50.37 0.76
TVAE+TCM 0.58 87.33 0.90 89.01 5.77 30.87 2.48 87.00 2.13 82.57 1.56 54.98 7.02 49.43 0.04
TVAE+TCMP 0.34 88.42 0.26 88.02 0.72 29.78 1.42 82.76 0.92 77.92 0.66 52.52 7.24 51.35 0.29
TVAE+DCM 0.45 88.12 0.32 87.54 1.24 31.54 0.38 83.45 0.84 78.57 0.68 54.85 5.05 48.81 2.10

### v2

@@ Abstract
Diffusion models have shown strong performance in generating high-quality tabular data, but they carry privacy risks by inadvertently reproducing exact training samples. While prior work focuses on dataset-level augmentation for memorization mitigation, little is known about which individual samples contribute most to memorization. In this paper, we present the first data-centric study of memorization dynamics in tabular diffusion models. We begin by quantifying memorization for each real sample based on how many generated samples are flagged as its memorized replicas, using a relative distance ratio metric.111We follow Yoon et al. (2023); Gu et al. (2023); Fang et al. (2024) to define sample memorization using the relative distance ratio. Our empirical analysis reveals a heavy-tailed distribution of memorization counts: a small subset of samples disproportionately contributes to leakage, a finding further validated through sample-removal experiments. To better understand this effect, we divide real samples into the top- and non-top-memorized groups (tags) and analyze their training-time behavior differences. We track when each sample is first memorized and monitor per-epoch memorization intensity (AUC) across groups. We find that memorized samples tend to be memorized slightly earlier and show significantly stronger memorization signals in early training stages. Based on these insights, we propose DynamicCut, a two-stage, model-agnostic mitigation method. DynamicCut (a) ranks real samples by their epoch-wise memorization intensity, (b) prunes a tunable top fraction, and (c) retrains the model on the filtered dataset. Across multiple benchmark tabular datasets and tabular diffusion models, DynamicCut reduces memorization ratios with negligible impact on data diversity and downstream task performance, and complements existing data augmentation methods for further memorization mitigation. Furthermore, DynamicCut has transferability across different generative models for memorization sample tagging, i.e., high-ranked samples identified from one model (e.g., a diffusion model) are also effective in reducing memorization when removed from other generative models such as GANs and VAEs.
@@ 1 Introduction
Despite the impressive performance of diffusion models in synthesizing high-quality tabular data, memorization remains a critical concern. In particular, tabular generative models risk reproducing individual training records that, while appearing structurally altered, still retain sensitive attributes of real individuals. This subtle form of privacy leakage violates core principles of generative modeling and has severe implications in sensitive domains such as healthcare and finance Karras et al. (2022); Carlini et al. (2021); Song et al. (2021); Ho et al. (2020). To address this, prior work has primarily focused on memorization detection based on nearest-neighbor distance ratios and dataset-level augmentation strategies such as feature mixing TabCutMix Fang et al. (2024). However, these methods fall short in two key ways. First, the memorization detection tool does not explain when or which samples are memorized during training. Second, the memorization mitigation method like TabCutMix operates uniformly across the dataset, ignoring the heterogeneous nature of sample-level memorization risks. This gap in understanding motivates a central question:
@@ 2 Related Work
Memorization has been widely studied in image and language generation van den Burg and Williams (2021); Gu et al. (2023); Somepalli et al. (2023b); Wen et al. (2024); Hintersdorf et al. (2024); Huang et al. (2024); Shah et al. (2025), where models often replicate training data. In vision, diffusion models like Stable Diffusion Rombach et al. (2022) and DDPM Ho et al. (2020) exhibit memorization Somepalli et al. (2023a); Carlini et al. (2021), mitigated by methods such as concept ablation Kumari et al. (2023) and adaptive sample suppression Chen et al. (2024). In language models, techniques like Goldfish Loss Hans et al. (2024) and early memorization prediction Biderman et al. (2024) aim to reduce overfitting to specific sequences. However, memorization in tabular data generation remains underexplored. Due to the structured and mixed-type nature of tabular data, understanding how memorization arises—and how it can be mitigated—presents a distinct and open research challenge.
@@ 4.1 Memorization and Forgetting: Event Definitions
We adopt the widely used relative distance ratio criterion Yoon et al. (2023); Gu et al. (2023); Fang et al. (2024) to determine whether a generated sample memorizes a real training sample. Formally, let be the training dataset and be a distance metric in the input space. Let and denote the nearest and second-nearest training samples to , respectively. We define the distance ratio as , A sample is considered memorized if , where is a fixed memorization threshold (we follow prior work and use ).
@@ 5.2 Algorithms
| Methods | Mem. Ratio (%) | Improv. | MLE (%) | -Precision(%) | -Recall(%) | Shape Score(%) | Trend Score(%) | C2ST(%) | DCR(%) | |
Default |
TabDDPM | 0.45 | - | 76.79 0.69 | 98.15 1.45 | 44.41 0.70 | 97.58 0.95 | 94.46 0.68 | 91.85 6.04 | 49.12 0.94 |
| TabDDPM+DC | 1.04 | 76.42 0.73 | 93.26 5.44 | 42.64 1.16 | 93.40 3.46 | 89.65 4.00 | 86.09 6.48 | 50.78 1.60 | ||
| CTGAN | 0.63 | - | 68.68 0.21 | 68.95 1.70 | 16.49 0.55 | 85.04 0.93 | 77.05 2.47 | 61.89 4.89 | 49.45 0.80 | |
| CTGAN+DC | 0.20 | 70.04 0.66 | 69.06 3.69 | 16.46 0.97 | 85.11 0.20 | 77.03 0.41 | 58.03 5.65 | 50.93 1.82 | ||
| TVAE | 0.43 | - | 72.24 0.36 | 82.97 0.37 | 20.57 0.42 | 89.37 0.54 | 83.36 0.99 | 52.63 0.37 | 51.33 0.96 | |
| TVAE+DC | 0.60 | 76.87 0.54 | 82.54 1.97 | 21.25 0.60 | 89.98 0.74 | 80.36 0.63 | 51.42 3.67 | 50.01 2.21 | ||
Shoppers |
TabDDPM | 0.31 | - | 92.17 0.32 | 93.16 1.58 | 52.57 1.30 | 97.08 0.46 | 92.92 3.27 | 86.74 0.63 | 51.36 0.63 |
| TabDDPM+DC | 2.38 | 91.08 1.18 | 91.88 2.74 | 52.74 2.67 | 93.56 2.89 | 89.40 3.87 | 80.12 4.54 | 51.87 0.23 | ||
| CTGAN | 1.15 | - | 83.22 1.30 | 85.26 5.17 | 26.67 1.66 | 77.73 0.50 | 86.48 0.69 | 67.18 5.84 | 48.54 0.74 | |
| CTGAN+DC | 0.45 | 83.71 0.44 | 89.13 4.17 | 27.73 0.66 | 77.78 0.87 | 86.93 0.34 | 69.31 6.11 | 48.55 1.50 | ||
| TVAE | 0.77 | - | 86.65 0.48 | 50.32 2.98 | 11.16 1.33 | 74.75 0.92 | 77.56 1.11 | 21.29 3.28 | 43.31 0.65 | |
| TVAE+DC | 0.83 | 86.93 0.38 | 50.61 1.39 | 12.49 0.85 | 74.66 0.73 | 79.14 0.30 | 21.38 2.28 | 47.61 5.38 | ||
Adult |
TabDDPM | 0.18 | - | 91.09 0.07 | 93.58 1.99 | 51.52 2.29 | 98.84 0.03 | 97.78 0.07 | 94.63 1.19 | 51.56 0.34 |
| TabDDPM+DC | 1.03 | 90.81 0.29 | 97.45 1.20 | 45.54 1.32 | 95.68 2.55 | 93.62 2.82 | 86.84 3.48 | 50.42 1.24 | ||
| CTGAN | 0.62 | - | 88.64 0.32 | 78.15 3.66 | 26.27 0.67 | 82.43 0.90 | 82.83 0.93 | 63.64 2.74 | 49.14 0.27 | |
| CTGAN+DC | 0.27 | 88.23 0.67 | 74.41 1.07 | 23.52 0.92 | 81.08 1.56 | 83.26 1.00 | 61.72 4.20 | 48.14 0.51 | ||
| TVAE | 0.41 | - | 88.61 0.49 | 92.34 2.19 | 29.90 1.00 | 82.71 0.45 | 79.06 0.90 | 48.67 2.70 | 48.76 0.25 | |
| TVAE+DC | 1.10 | 88.25 0.33 | 87.80 0.74 | 30.57 2.82 | 83.54 0.56 | 78.71 0.86 | 55.64 2.59 | 45.96 0.54 |
@@ 6.2 Unified Evaluation of Memorization under Data Quality and Training Dynamics
| Methods | Mem. Ratio (%) | Improv. | MLE (%) | -Precision(%) | -Recall(%) | Shape Score(%) | Trend Score(%) | C2ST(%) | DCR(%) | |
Default |
TabDDPM | 0.45 | - | 76.79 0.69 | 98.15 1.45 | 44.41 0.70 | 97.58 0.95 | 94.46 0.68 | 91.85 6.04 | 49.12 0.94 |
| TabDDPM+SMOTE | 0.51 | 76.92 0.35 | 91.19 0.68 | 40.52 0.65 | 94.89 1.46 | 28.63 2.28 | 72.73 0.69 | 50.95 0.38 | ||
| TabDDPM+TCM | 0.47 | 76.47 0.60 | 97.30 0.46 | 38.72 2.78 | 97.27 1.74 | 93.27 2.52 | 94.72 3.87 | 50.23 0.53 | ||
| TabDDPM+TCMP | 0.24 | 76.92 0.17 | 98.26 0.25 | 41.92 0.52 | 97.37 0.09 | 91.42 1.15 | 95.64 0.49 | 49.75 0.32 | ||
| TabDDPM+DCM | 2.06 | 76.68 0.76 | 91.97 1.78 | 37.89 0.84 | 93.62 2.78 | 89.30 3.02 | 87.63 2.15 | 48.17 2.59 | ||
| CTGAN | 0.63 | - | 68.68 0.21 | 68.95 1.70 | 16.49 0.55 | 85.04 0.93 | 77.05 2.47 | 61.89 4.89 | 49.45 0.80 | |
| CTGAN+SMOTE | 0.41 | 71.23 1.54 | 65.18 2.81 | 17.13 0.20 | 84.31 0.20 | 31.76 0.11 | 56.35 0.60 | 49.05 2.85 | ||
| CTGAN+TCM | 0.28 | 73.21 0.45 | 75.30 2.12 | 16.63 0.27 | 85.77 0.13 | 76.69 1.20 | 58.62 5.34 | 49.90 3.12 | ||
| CTGAN+TCMP | 0.25 | 70.05 0.81 | 71.33 0.82 | 17.02 0.60 | 85.28 0.97 | 78.09 0.32 | 63.67 3.07 | 50.14 0.60 | ||
| CTGAN+DCM | 0.07 | 72.37 0.39 | 71.72 0.74 | 16.64 0.20 | 85.78 0.70 | 76.85 1.62 | 57.98 3.09 | 48.48 0.24 | ||
| TVAE | 0.43 | - | 72.24 0.36 | 82.97 0.37 | 20.57 0.42 | 89.37 0.54 | 83.36 0.99 | 52.63 0.37 | 51.33 0.96 | |
| TVAE+SMOTE | 0.41 | 75.41 0.42 | 82.47 0.42 | 20.77 0.73 | 88.79 0.08 | 79.34 2.40 | 53.28 3.26 | 49.20 1.54 | ||
| TVAE+TCM | 0.37 | 75.05 0.66 | 80.23 2.07 | 21.90 2.36 | 89.65 0.15 | 80.15 0.75 | 52.10 4.29 | 49.99 1.78 | ||
| TVAE+TCMP | 0.28 | 72.75 0.52 | 81.57 0.32 | 19.52 0.39 | 89.33 0.48 | 78.04 6.05 | 45.60 0.90 | 50.07 0.52 | ||
| TVAE+DCM | 0.21 | 76.27 0.25 | 85.49 0.41 | 21.66 0.19 | 90.36 0.25 | 86.76 5.13 | 54.81 0.79 | 51.07 1.35 | ||
Shoppers |
TabDDPM | 0.31 | - | 92.17 0.32 | 93.16 1.58 | 52.57 1.30 | 97.08 0.46 | 92.92 3.27 | 86.74 0.63 | 51.36 0.63 |
| TabDDPM+SMOTE | 1.46 | 89.96 0.95 | 94.41 4.67 | 45.22 3.26 | 90.78 0.49 | 83.09 2.47 | 64.05 1.44 | 51.94 1.52 | ||
| TabDDPM+TCM | 1.17 | 92.17 0.26 | 94.41 1.49 | 50.05 1.59 | 97.18 0.34 | 93.95 0.51 | 86.96 0.50 | 47.52 1.81 | ||
| TabDDPM+TCMP | 0.35 | 92.09 0.99 | 93.43 1.65 | 52.30 0.73 | 97.31 0.22 | 94.79 0.30 | 87.02 2.04 | 50.83 0.59 | ||
| TabDDPM+DCM | 2.83 | 91.46 0.74 | 93.25 3.87 | 46.15 4.42 | 92.32 5.00 | 89.11 4.95 | 80.56 4.81 | 49.74 3.34 | ||
| CTGAN | 1.15 | - | 83.22 1.30 | 85.26 5.17 | 26.67 1.66 | 77.73 0.50 | 86.48 0.69 | 67.18 5.84 | 48.54 0.74 | |
| CTGAN+SMOTE | 1.86 | 87.09 1.22 | 76.13 13.12 | 26.19 3.66 | 75.57 2.48 | 83.94 1.13 | 59.83 6.67 | 50.74 1.75 | ||
| CTGAN+TCM | 0.42 | 84.21 1.21 | 80.82 7.68 | 26.50 1.50 | 77.51 2.15 | 85.75 1.31 | 62.04 6.75 | 49.90 2.33 | ||
| CTGAN+TCMP | 0.35 | 83.94 0.80 | 82.07 9.62 | 23.47 0.43 | 76.15 0.22 | 84.56 0.40 | 64.60 0.82 | 52.02 0.13 | ||
| CTGAN+DCM | 0.56 | 85.51 0.46 | 83.51 1.60 | 26.19 0.82 | 78.23 0.30 | 86.36 0.47 | 68.85 2.45 | 50.64 1.11 | ||
| TVAE | 0.77 | - | 86.65 0.48 | 50.32 2.98 | 11.16 1.33 | 74.75 0.92 | 77.56 1.11 | 21.29 3.28 | 43.31 0.65 | |
| TVAE+SMOTE | 0.39 | 89.71 0.69 | 52.39 3.94 | 13.74 0.54 | 71.80 0.63 | 76.20 0.68 | 20.67 2.34 | 47.42 3.40 | ||
| TVAE+TCM | 2.38 | 87.33 0.90 | 49.75 5.04 | 12.66 3.18 | 75.47 1.30 | 78.91 1.30 | 21.87 3.87 | 50.06 1.78 | ||
| TVAE+TCMP | 0.80 | 87.07 0.74 | 50.69 0.44 | 10.54 1.68 | 74.86 0.26 | 78.55 0.48 | 21.78 4.28 | 43.21 0.69 | ||
| TVAE+DCM | 1.33 | 88.32 0.29 | 51.12 2.46 | 11.98 0.58 | 75.02 0.97 | 79.02 0.12 | 22.99 2.12 | 49.98 2.55 | ||
Adult |
TabDDPM | 0.18 | - | 91.09 0.07 | 93.58 1.99 | 51.52 2.29 | 98.84 0.03 | 97.78 0.07 | 94.63 1.19 | 51.56 0.34 |
| TabDDPM+SMOTE | 0.78 | 90.41 0.36 | 94.93 1.72 | 46.10 0.65 | 93.40 1.12 | 90.76 1.76 | 80.75 0.84 | 51.82 0.56 | ||
| TabDDPM+TCM | 0.19 | 91.15 0.06 | 94.97 0.06 | 47.43 1.46 | 98.65 0.03 | 97.75 0.07 | 85.61 16.03 | 50.99 0.65 | ||
| TabDDPM+TCMP | 2.11 | 90.54 0.17 | 92.26 6.97 | 43.49 3.74 | 95.10 4.27 | 91.50 6.53 | 84.76 10.12 | 50.68 0.89 | ||
| TabDDPM+DCM | 0.57 | 90.81 0.16 | 95.23 1.14 | 45.76 1.02 | 98.27 0.32 | 96.71 0.86 | 92.64 1.65 | 50.79 0.46 | ||
| CTGAN | 0.62 | - | 88.64 0.32 | 78.15 3.66 | 26.27 0.67 | 82.43 0.90 | 82.83 0.93 | 63.64 2.74 | 49.14 0.27 | |
| CTGAN+SMOTE | 0.69 | 88.59 0.41 | 74.96 5.01 | 24.65 1.44 | 82.16 1.14 | 80.23 0.38 | 60.01 2.10 | 51.71 2.15 | ||
| CTGAN+TCM | 0.48 | 88.94 0.64 | 76.10 4.79 | 26.29 1.53 | 83.26 1.30 | 80.09 1.89 | 61.16 3.59 | 49.87 0.64 | ||
| CTGAN+TCMP | 0.31 | 88.63 0.66 | 76.27 0.60 | 25.56 0.33 | 81.42 0.32 | 82.11 1.01 | 60.98 1.67 | 50.96 0.17 | ||
| CTGAN+DCM | 0.88 | 88.45 0.29 | 79.31 1.04 | 23.31 1.42 | 82.23 0.23 | 82.80 2.27 | 63.46 3.43 | 48.13 0.33 | ||
| TVAE | 0.41 | - | 88.61 0.49 | 92.34 2.19 | 29.90 1.00 | 82.71 0.45 | 79.06 0.90 | 48.67 2.70 | 48.76 0.25 | |
| TVAE+SMOTE | 0.16 | 87.89 0.32 | 89.00 3.48 | 31.01 2.50 | 80.65 1.28 | 81.83 1.64 | 52.98 6.01 | 50.37 0.76 | ||
| TVAE+TCM | 0.58 | 87.33 0.90 | 89.01 5.77 | 30.87 2.48 | 87.00 2.13 | 82.57 1.56 | 54.98 7.02 | 49.43 0.04 | ||
| TVAE+TCMP | 0.34 | 88.42 0.26 | 88.02 0.72 | 29.78 1.42 | 82.76 0.92 | 77.92 0.66 | 52.52 7.24 | 51.35 0.29 | ||
| TVAE+DCM | 0.45 | 88.12 0.32 | 87.54 1.24 | 31.54 0.38 | 83.45 0.84 | 78.57 0.68 | 54.85 5.05 | 48.81 2.10 |